HUBBLE SPACE TELESCOPE DAILY REPORT #999 PERIOD COVERED: 0800 EST Daily Status Report as of 295/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 5075 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 1) Observations were made to make a standard calibration to obtain Earth flats using the Wide Field Camera. A total of two observations were taken. One was reported to be saturated, and one was okay. 1.2 Completed PC 4308 (Snapshot Survey for Companions to Planetary Nebula Nuclei - Cycle 3 High) The Planetary Camera was used to make an observation of the central region of planetary nebula ABELL64. It is possible that the central star in a double star. Two observations was received and displayed in the OSS, the data was reported to be normal. 1.3 Completed PC 4904 (The Nuclei of Nearby S0 and Spiral Galaxies - III A PC Snapshot Images) The Planetary Camera was used to make one observations of the low-red-shift galaxy NGC-1574 to study its surface brightness profile, identify sub-arc-second scale condensations, find point sources, and look for gasious disk and dust lanes. The observation was executed without problems. The data was received and displayed. No problems were reported. 1.4 Completed FOS/BL 4260 (Extreme Stellar Winds and Post Main Sequence Evolution) The Faint Object Spectrograph was used to make observations of star BE261. This star is believed to be intermediate between a class O and WR star and to have strong winds and high mass loss rate (i.e., much higher than normal O class stars). The onboard target acquisition was successful. The peakup had its greatest signal in the final group, so it could not be judged if the peakup failed or succeeded. The spectral observations were received and displayed. There was some uncertainty about the target being in the slit. No other anomalies were reported. 1.5 Completed WFC 3589 (Parallel High Resolution Imaging of Defuse Objects in the Magellanic Clouds) The Wide Field Planetary Camera was used to make observations of the nebulous object LMC-N44D which is located in the Large Magellanic Cloud. This observations was take in parallel with the prime science observation. Am emission line filter and nearby continuum filter were used. One image was received and displayed. It was reported to look okay. 1.6 Completed HSP/UV1/POL 4209 (HSP Detector Sensitivity Test) The High Speed Photometer was used to make four calibration observations of standard star HD+75D325 to determine the instruments long term sensitivity of the IDT detectors. The observations could not be monitored because of hardware problems in the OSS. 1.7 Completed HSP/UV1 4945 (Periodic Variations in DQ Herculis Stars - Cycle 2 High) The High Speed Photometer was used to make nearly five hours of observations of variable star DQ Her (aka, Nova Her 1934). The chosen bandpass avoids the strong ultraviolet emission seen in the spectrum of this target. Rapid periodic luminous variations in the optical and X-Ray are characteristic for this class of cataclysmic variable. The observations could not be monitored because of hardware problems in the OSS. 1.8 Completed GHRS/2 5040 (GHRS Detector Calibrations - Cycle 3) The High Resolution spectrograph was used to make a set of internal calibrations to check the stability of the HRS detectors. The test included dark counts, focus checks. and flat field observations. The observation was executed as scheduled. The observations could not be monitored because of hardware problems in the OSS. 1.9 Completed HSP/UV2 4801 (Quasi-Periodic Oscillations in AM Herculis Binaries) The High speed Photometer was used to make ultraviolet high speed observations of the oscillations in the light from the variable star AM Her. This is replacement of proposal 3607. This binary system has a magnetic white dwarf star which is accreting matter from its companion. The observation was executed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Significant HSTARs: None 2.4 Operations Note: The SMGT#9 RGA test successfully completed yesterday, Thursday Oct. 21, 1993 from 8am until 8pm. All simulation objectives were met, and no problems have been reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Four Sets of PC 4308 (Snapshot Survey for Companions to Planetary Nebula Nuclei - Cycle 3 High) 3.2 Complete GHRS/2 4573 (Boron in Orion: Cycle 3 Medium) 3.3 Complete WFC 5025 (WF/PC Cycle Calibration: Photometric Calibration Monitor) 3.4 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.5 Complete Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.6 Complete FOS BL 4260 (Extreme Stellar Winds and Post Main Sequence Evolution) 3.7 Complete Two Set of PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.8 Complete WFC 3589 (Parallel High Resolution Imaging of Defuse Objects in the Magellanic Clouds) 3.9 Complete PC 4940 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) 3.10 Complete FGS 5118 (Monitor the Long Term Stability of FGS Revised) 3.11 Complete FOS/RD 4784 (Transition Comet: Search for OH) 3.12 Complete Two Sets of PC 4903 (The Nuclei of Nearby S0 and Spiral Galaxies - II A PC Snapshot Imaging) 3.13 Complete FOS/RD 4581 (Quasar Absorption Line Survey - Cycle 3 High) 3.14 Complete Two Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.15 Complete FOS/BL 2245 (Ultraviolet Interstellar Polarization) 3.16 Complete WFC 5076 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 2) 3.17 Complete FOS/BL 5078 (Spectroscopic Observations of the Exposed White Dwarfs) 3.18 Complete FOC/96 5028 (FOC Objective Prisms - Wavelength Calibration Verification) 3.19 Complete FGS 4893 (Orbital Motions of Close Binary Stars - cycle 3 High) 3.20 Complete WFC 4849 (GTO High Latitude Program - Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-